solve by addition
x - y = -3
x - y = 4

Answer by wuwei96815(245)

This set of equations does not have a solution.

Want proof?

x-y=-3 times -1 yields -x+y=3

Now add the two equations together:
-x+y=3 plus x-y=4 yields 0=7

Do you see now why there is no solution?